Suppose the market for grass seed can be expressed as follows.

Demand: QD = 100 - 2p

Supply: QS = 3p

At the market equilibrium, calculate the price elasticities of supply and demand? What would happen with the quantity demanded if price reaches to 30? Price reaches to 50?


1
Expert's answer
2020-11-09T10:35:38-0500
"100-2p=3p"

"p=20"

"Q=60"


"E_D=-2\times\frac{20}{60}=-0.67"


"E_S=3\times \frac {20}{60}=1"

If p=30


"Q=3\times30-(100-2\times30)=90-40=50"

There will be a shortage of 50 items


If p=50


"Q=3\times50-(100-20\times 50)=150"

There will be an excess of 150 pcs.